OpenStudy (anonymous):

Find dxy if y=e^(1/x^2) + (1/e^x^2)

zepdrix (zepdrix):

what is \(dxy\)? Is that suppose to be \(\dfrac{dy}{dx}\)?

OpenStudy (anonymous):

Dy (with a tiny x in the middle of D and y). Im assuming its implicit differentiation

OpenStudy (anonymous):

just use the product rule since both x^2 and e^x are functions of x. the product rule states that dy/dx is the derivative of the first function times the second plus the first function times the derivative of the second. so dy/dx = dydx(x^2) * (e^x) + (x^2) * dy/dx(e^x) dy/dx (x^2) is 2x dy/dx (e^x) is e^x so you get: 2x* e^x + x^2 * e^x now you can factor out x*e^x from both terms to get xe^x(2+x)

zepdrix (zepdrix):

\[\huge y=e^{\left(1/x^2\right)}+\dfrac{1}{e^{x^2}}\] Hmm I don't see product rule anywhere. Are you sure you read the problem correctly @jk3012 ? I know these can be tricky to read when they're not formatted :C

zepdrix (zepdrix):

It looks like we're actually going to be applying the chain rule several times :O Let's clean up the exponents before differentiating.\[\huge y=e^{\left(x^{-2}\right)}+e^{-x^2}\]Understand how I changed those? I took the terms out of the denominator and applied a negative power to them.

OpenStudy (anonymous):

ah ok, i had originally gotten the conversion to negative power wrong

zepdrix (zepdrix):

Maybe it makes more sense to write it like this.\[\huge y=e^{(x^{-2})}+e^{-\left(x^2\right)}\]The negative is being applied to the exponent on the X in the first term, and to the exponent on the \(e\) in the second term.

zepdrix (zepdrix):

Need to see some steps after that?

OpenStudy (anonymous):

yes please...

OpenStudy (anonymous):

would it be [e^(x^-2)](-2x^-3) + [e^-(x^2)](-2x)?

zepdrix (zepdrix):

\[\huge y'=e^{(x^{-2})}\color{royalblue}{(x^{-2})'}+e^{-\left(x^2\right)}\color{royalblue}{(-x^2)'}\] Yesss very good! :)

OpenStudy (anonymous):

and thats it right?

OpenStudy (anonymous):

is there a need to simplify even more?

zepdrix (zepdrix):

It looks a little nicer if you pull the x's to the front, but yes that's all! :)

zepdrix (zepdrix):

And pull the negatives to the fronts also *

OpenStudy (anonymous):

ok will do... thanks so much :)